The Central Intelligence Agency (CIA) has specialists who analyze the frequencies of letters of the alphabet in an attempt to decipher intercepted messages. In standard English text, the letter t is used at a rate of 6.95%.

Find the mean and standard deviation for the number of times the letter t will be found on atypical page of 2500 characters. (Round answer to one decimal place.)

μ= 173.8 (Round answer to two decimal places.)

σ= 12.72

Give the interval of usual values for the number of t's appearing on a page of 2500 characters. (Enter answer as an interval using square-brackets only with whole numbers.) usual values =
Answer #1

Usual values:

Min = μ-2σ = 173.8-2*12.72 = 148.36 = 148

Max = μ+2σ = 173.8+2*12.72 = 199.24 = 199

Ans : (148, 199)